# Subsystem fidelity in two-dimensional conformal field theories

### v1

@@ 1 Introduction
In this paper, we focus on calculating subsystem fidelity in two-dimensional conformal field theory (2D CFT). Fidelity measures the similarity between two quantum states. For two density matrices $\rho$ and $\sigma$, it is defined as [1, 2]
$F(\rho,\sigma)=\textrm{tr}\sqrt{\sqrt{\rho},\sigma\sqrt{\rho}}.$ (1.1)
@@
$F(\rho,\sigma)=\lim_{p\to 1/2}\textrm{tr}[(\rho^{\frac{1-p}{2p}}\sigma\rho^{\frac{1-p}{2p}})^{p}].$ (1.2)
@@
More recently, a refined replica trick was introduced and applied in the context of AdS/CFT correspondence in [21]
@@
We study subsystem fidelity in general 2D CFTs using twist operators [19, 22, 23] and their operator product expansion (OPE) [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 9, 10]. By evaluating the short-interval expansion of subsystem fidelity via the OPE of twist operators, we show that it receives contributions from various quasiprimary operators ordered by their scaling dimensions. We derive universal contributions from general quasiprimary operators in arbitrary 2D CFTs, as well as specific contributions from particular operators in free boson and free fermion theories. Our results in these solvable models agree with existing analytical and numerical calculations. Furthermore, we extend the method to 2D holographic CFTs, where through the AdS3/CFT2 correspondence [35], we analyze the perturbative distinguishability of black hole microstates.
@@ 2.2 OPE of twist operators
We define the “generalized fidelity” as
@@
and obtain fidelity from the limit [21]
@@
$\langle\mathcal{T}(\ell,\ell)\tilde{\mathcal{T}}(0,0)\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{2(h_{n}+\bar{h}_{n})}\Big(1+\sum_{K}\ell^{\Delta_{K}}d_{K}\langle\Phi_{K}\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}\Big).$ (2.15)
@@
$F_{p}(\rho_{A},\sigma_{A})=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{2(h_{n}+\bar{h}_{n})}\bigg(1+\sum_{k=1}^{n}\sum_{\{\mathcal{X}_{1},\cdots,\mathcal{X}_{k}\}}\ell^{\Delta_{\mathcal{X}_{1}}+\cdots+\Delta_{\mathcal{X}_{k}}}\mathcal{F}_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{(p)}\bigg),$ (2.18)
@@ 2.3 Universal contributions from $\mathcal{X}\mathcal{X}$
with shorthand $j_{12}\equiv j_{1}-j_{2}$. The O(n1)O(n-1) part exists if $\mathcal{X}$ is quasiprimary but not primary. Since we take the $n\to 1$ limit for the final result, the O(n1)O(n-1) part does not contribute to fidelity. The contribution from $\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}$ is the $p\to\frac{1}{2}$ limit (i.e., $n\to 1$ limit) of
@@
where $j_{12}\equiv j_{1}-j_{2}$.
To evaluate this sum, we separate the replica indices into even and odd values. Denoting even integers $a=0,2,\cdots,2p-2$ and odd integers $b=1,3,\cdots,2p-1$, we get
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}}=(\langle\mathcal{X}\rangle\rho^{2}+\langle\mathcal{X}\rangle_{\sigma}^{2})\sum_{a_{1}<a_{2}}d_{\mathcal{X}\mathcal{X}}^{a_{1}a_{2}}+\langle\mathcal{X}\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}\sum_{a,b}d_{\mathcal{X}\mathcal{X}}^{ab}.$ (2.24)
@@ 2.4 Universal contributions from $\mathcal{X}\mathcal{X}\mathcal{X}$
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=(\langle\mathcal{X}\rangle\rho^{3}+\langle\mathcal{X}\rangle_{\sigma}^{3})\sum_{a_{1}<a_{2}<a_{3}}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}a_{3}}+(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})\langle\mathcal{X}\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}\sum_{a_{1}<a_{2},b}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}b}.$ (2.30)
@@
$\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=\frac{\Gamma(\Delta\mathcal{X}+\frac{1}{2})^{2}}{2^{2\Delta_{\mathcal{X}}+4}\pi\Gamma(\frac{\Delta_{\mathcal{X}}}{2}+1)\Gamma(\frac{3\Delta_{\mathcal{X}}}{2}+1)}\frac{\mathrm{i}^{s_{\mathcal{X}}}C_{\mathcal{X}\mathcal{X}\mathcal{X}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})}{\alpha_{\mathcal{X}}^{3}}.$ (2.33)
@@ 2.5 Universal contributions from $T\mathcal{X}\mathcal{X}$
$\mathcal{F}_{\bar{T}\mathcal{X}\mathcal{X}}^{(1/2)}=\frac{\bar{h}_{\mathcal{X}}\Gamma(\Delta_{\mathcal{X}}+\frac{1}{2})}{2^{2\Delta_{\mathcal{X}}+5}\sqrt{\pi}c\Gamma(\Delta_{\mathcal{X}}+2)}\frac{\mathrm{i}^{2s_{\mathcal{X}}}}{\alpha_{\mathcal{X}}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})\big[(\langle\bar{T}\rangle_{\rho}+\langle\bar{T}\rangle_{\sigma})(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})$ (2.37)
@@ 2.6 Contributions from $JJJJ$ in free massless boson theory
In free massless boson theory, there is the the current operator $J$, which is primary and has normalization $\alpha_{J}=1$ and conformal weights $(1,0)$. The OPE coefficient of CFTn primary operator $J_{j_{1}}J_{j_{2}}J_{j_{3}}J_{j_{4}}$ with $0\leq j_{1}<j_{2}<j_{3}<j_{4}\leq n-1$ is [42]
@@
where $G_{6}(p)$, $G_{7}(p)$, and $G_{8}(p)$ are defined in (A.18), (A.19), and (A), respectively. Using the analytical continuations (A.20), we obtain contributions from JJJJJJJJ
@@ 3.1 Case I
Using the universal contributions derived in the previous section, specifically equations (2.27) and (2.33), and the factorization formula
$F(\rho_{A},\sigma_{A})=\big[1+\ell^{4}\mathcal{F}^{(1/2)}{TT}+\ell^{6}\mathcal{F}^{(1/2)}{TTT}+O(\ell^{8})\big]\big[1+\ell^{4}\mathcal{F}^{(1/2)}{\bar{T}\bar{T}}+\ell^{6}\mathcal{F}^{(1/2)}{\bar{T}\bar{T}\bar{T}}+O(\ell^{8})\big]$
@@ 3.2 Case II
Using the contributions from current operators (2.27), (2.33), (2.40), and (2.41), and organizing them in a factorized form:
$F(\rho_{A},\sigma_{A})=\big[1+\ell^{2}\mathcal{F}^{(1/2)}{JJ}+\ell^{4}\big(\mathcal{F}^{(1/2)}{JJJJ}+\mathcal{F}^{(1/2)}{TJJ}+\mathcal{F}^{(1/2)}{TT}\big)+O(\ell^{6})\big]$
@@
we reproduce the short interval expansion of the vertex operator fidelity. The factorization reflects the holomorphic-antiholomorphic separation in the free boson theory.

### v2

@@ 1 Introduction
In this paper, we focus on calculating subsystem fidelity in two-dimensional conformal field theories (2D CFTs). Fidelity measures the similarity between two quantum states. For two density matrices $\rho$ and $\sigma$, it is defined as [1, 2]
$F(\rho,\sigma)=\textrm{tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}.$ (1.1)
@@
$F(\rho,\sigma)=\lim_{p\to 1/2}\textrm{tr}\big[\big(\rho^{\frac{1-p}{2p}}\sigma\rho^{\frac{1-p}{2p}}\big)^{p}\big].$ (1.2)
@@
More recently, a refined replica trick was introduced [21, 22]
@@
Although not immediately apparent, the new definition of fidelity (1.4), proposed in [20], is in fact equivalent to the conventional expression (1.1). This equivalence can be understood via the replica trick. To evaluate the standard fidelity (1.1), one employs the identity
$F(\rho,\sigma)=\lim_{p\to 1/2}\textrm{tr}\big[(\sqrt{\rho}\sigma\sqrt{\rho})^{p}\big].$ (1.6)
Meanwhile, for any nonnegative integer $p$, the identity $\textrm{tr}[(\sqrt{\rho}\sigma\sqrt{\rho})^{p}]=\textrm{tr}[(\rho\sigma)^{p}]$ always holds. As a result, the replica trick in (1.5) yields the same limit as (1.6), confirming that both definitions coincide. Alternatively, one may observe that for any matrices $X$ and $Y$ of compatible dimensions and any nonnegative integer $p$, the equality $\textrm{tr}[(XY)^{p}]=\textrm{tr}[(YX)^{p}]$ implies that $XY$ and $YX$ share the same nonvanishing eigenvalues. In particular, $\sqrt{X}Y\sqrt{X}$ and $XY$ are isospectral, provided the relevant matrix square roots exist. Consequently, their square roots, $\sqrt{\sqrt{X}Y\sqrt{X}}$ and $\sqrt{XY}$, also have identical eigenvalues. This reasoning again shows that definitions (1.1) and (1.4) are equivalent, regardless of whether $\rho$ and $\sigma$ commute.
We study subsystem fidelity in general 2D CFTs using twist operators [19, 23, 24] and their operator product expansion (OPE) [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 9, 10]. By evaluating the short-interval expansion of subsystem fidelity via the OPE of twist operators, we show that it receives contributions from various quasiprimary operators ordered by their scaling dimensions. We derive universal contributions from general quasiprimary operators in arbitrary 2D CFTs, as well as specific contributions from particular operators in free boson and free fermion theories. Our results in these solvable models agree with existing analytical and numerical calculations. Furthermore, we extend the method to 2D holographic CFTs, where through the AdS3/CFT2 correspondence [37], we analyze the perturbative distinguishability of black hole microstates.
@@ 2.2 OPE of twist operators
We define the “generalized fidelity” as [^2] The “Rényi fidelity” was introduced in [21] as $F_{p}(\rho,\sigma)=\frac{\textrm{tr}[(\rho\sigma)^{p}]}{\sqrt{\textrm{tr}(\rho^{2p})\textrm{tr}(\rho^{2p})}}.$ The special $p=1$ case of the Rényi fidelity was proposed in [41] and has since been applied to problems in thermalization and revival in two-dimensional conformal field theories [5], bulk reconstruction in AdS/CFT correspondence [17], and non-equilibrium evolution after a quantum quench [42, 43].
@@
and obtain fidelity from the limit [21, 22]
@@
$\langle\mathcal{T}(\ell,\ell)\tilde{\mathcal{T}}(0,0)\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{-2(h_{n}+\bar{h}_{n})}\Big(1+\sum_{K}\ell^{\Delta_{K}}d_{K}\langle\Phi_{K}\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}\Big).$ (2.15)
@@
$F_{p}(\rho_{A},\sigma_{A})=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{-2(h_{n}+\bar{h}_{n})}\bigg(1+\sum_{k=1}^{n}\sum_{\{\mathcal{X}_{1},\cdots,\mathcal{X}_{k}\}}\ell^{\Delta_{\mathcal{X}_{1}}+\cdots+\Delta_{\mathcal{X}_{k}}}\mathcal{F}_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{(p)}\bigg),$ (2.18)
@@
Formula (2.15) applies only to translationally invariant states; generalization to inhomogeneous states is possible but more complicated. As with entanglement entropy, fidelity in such cases depends on position of the subsystem. From the OPE of twist operators we have [28, 29]
$\langle\mathcal{T}(\ell,\ell)\tilde{\mathcal{T}}(0,0)\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{-2(h_{n}+\bar{h}_{n})}\Big(1+\sum_{K}\sum_{r,s=0}^{+\infty}\frac{a_{K}^{r}}{r!}\frac{\bar{a}_{K}^{s}}{s!}\ell^{\Delta_{K}+r+s}d_{K}\langle\partial^{r}\bar{\partial}^{s}\Phi_{K}(0,0)\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}\Big),$ (2.22)
where the coefficients are expressed through binomial coefficients as
$a_{K}^{r}\equiv\frac{C_{h_{K}+r-1}^{r}}{C_{2h_{K}+r-1}^{r}},~~\bar{a}_{K}^{s}\equiv\frac{C_{\bar{h}_{K}+s-1}^{s}}{C_{2\bar{h}_{K}+s-1}^{s}}.$ (2.23)
Moreover, the CFTn quasiprimary operators $\Phi_{K}$ are not restricted to the form (2.17); more general terms with derivatives must be included, such as [44]
$\mathcal{X}_{j_{1}}\mathrm{i}\partial\mathcal{X}_{j_{2}}-\mathrm{i}\partial\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}},$
$\mathcal{X}_{j_{1}}\partial\bar{\partial}\mathcal{X}_{j_{2}}+\partial\bar{\partial}\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}-\partial\mathcal{X}_{j_{1}}\bar{\partial}\mathcal{X}_{j_{2}}-\bar{\partial}\mathcal{X}_{j_{1}}\partial\mathcal{X}_{j_{2}},$
$\partial\mathcal{X}_{j_{1}}\partial\mathcal{X}_{j_{2}}-\frac{h_{\mathcal{X}}}{2h_{\mathcal{X}}+1}\left(\mathcal{X}_{j_{1}}\partial^{2}\mathcal{X}_{j_{2}}+\partial^{2}\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}\right).$ (2.24)
Interestingly, even for inhomogeneous states, the terms with derivatives do not modify the leading order in the short-interval expansion, and the result from subsection 2.3 remain dominant.
@@ 2.3 Universal contributions from $\mathcal{X}\mathcal{X}$
with shorthand $j_{12}\equiv j_{1}-j_{2}$. The contribution from $\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}$ is the $p\to\frac{1}{2}$ limit (i.e., $n\to 1$ limit) of
@@
where $j_{12}\equiv j_{1}-j_{2}$. For non-primary quasiprimary operators $\mathcal{X}$, the $\cdots$ terms in the conformal transformation (2.2) produce O(n1)O(n-1) contributions to $d_{\mathcal{X}\mathcal{X}}^{j_{1}j_{2}}$ in (2.25). These contributions remain of order O(n1)O(n-1) after summing over replica indices $j_{1},j_{2}$, vanish in the $n\to 1$ limit, and thus do not affect fidelity.
To evaluate the sum in (2.26), we separate the replica indices into even and odd values. Denoting even integers $a=0,2,\cdots,2p-2$ and odd integers $b=1,3,\cdots,2p-1$, we get
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}}=(\langle\mathcal{X}\rangle_{\rho}^{2}+\langle\mathcal{X}\rangle_{\sigma}^{2})\sum_{a_{1}<a_{2}}d_{\mathcal{X}\mathcal{X}}^{a_{1}a_{2}}+\langle\mathcal{X}\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}\sum_{a,b}d_{\mathcal{X}\mathcal{X}}^{ab}.$ (2.27)
@@ 2.4 Universal contributions from $\mathcal{X}\mathcal{X}\mathcal{X}$
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=(\langle\mathcal{X}\rangle_{\rho}^{3}+\langle\mathcal{X}\rangle_{\sigma}^{3})\sum_{a_{1}<a_{2}<a_{3}}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}a_{3}}+(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})\langle\mathcal{X}\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}\sum_{a_{1}<a_{2},b}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}b}.$ (2.33)
@@
$\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=\frac{\Gamma(\Delta_{\mathcal{X}}+\frac{1}{2})^{2}}{2^{2\Delta_{\mathcal{X}}+4}\pi\Gamma(\frac{\Delta_{\mathcal{X}}}{2}+1)\Gamma(\frac{3\Delta_{\mathcal{X}}}{2}+1)}\frac{\mathrm{i}^{s_{\mathcal{X}}}C_{\mathcal{X}\mathcal{X}\mathcal{X}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})}{\alpha_{\mathcal{X}}^{3}}.$ (2.36)
@@ 2.5 Universal contributions from $T\mathcal{X}\mathcal{X}$
$\mathcal{F}_{\bar{T}\mathcal{X}\mathcal{X}}^{(1/2)}=\frac{\bar{h}_{\mathcal{X}}\Gamma(\Delta_{\mathcal{X}}+\frac{1}{2})}{2^{2\Delta_{\mathcal{X}}+5}\sqrt{\pi}c\Gamma(\Delta_{\mathcal{X}}+2)}\frac{\mathrm{i}^{2s_{\mathcal{X}}}}{\alpha_{\mathcal{X}}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})\big[(\langle\bar{T}\rangle_{\rho}+\langle\bar{T}\rangle_{\sigma})(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})$
@@ 2.6 Contributions from $JJJJ$ in free massless boson theory
In free massless boson theory, there is the current operator $J$, which is primary and has normalization $\alpha_{J}=1$ and conformal weights $(1,0)$. The OPE coefficient of CFTn primary operator $J_{j_{1}}J_{j_{2}}J_{j_{3}}J_{j_{4}}$ with $0\leq j_{1}<j_{2}<j_{3}<j_{4}\leq n-1$ is [47]
@@
where $G_{6}(p)$, $G_{7}(p)$, and $G_{8}(p)$ are defined in (A.24), (A.25), and (A.26), respectively. Using the analytical continuations (A.27), we obtain contributions from JJJJJJJJ
@@ 3.1 Case I
Using the universal contributions derived in the previous section, specifically $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}}$ (2.30) with $\mathcal{X}=T,\bar{T}$ and $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}\mathcal{X}}$ (2.36) with $\mathcal{X}=T,\bar{T}$, and the factorization formula
$F(\rho_{A},\sigma_{A})=\big[1+\ell^{4}\mathcal{F}^{(1/2)}_{TT}+\ell^{6}\mathcal{F}^{(1/2)}_{TTT}+O(\ell^{8})\big]\big[1+\ell^{4}\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}}+\ell^{6}\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}\bar{T}}+O(\ell^{8})\big]$
@@
To clarify which operators contribute at which order of the short interval expansion for various fidelities, we list the nontrivial leading and subleading terms in Table 1 for both the free massless boson theory (discussed in this section) and the free massless fermion theory (discussed in the next section).
fidelity leading order subleading order
subsections 3.1 and 4.1 $\ell^{4}\big(\mathcal{F}^{(1/2)}_{TT}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}}\big)$ $\ell^{6}\big(\mathcal{F}^{(1/2)}_{TTT}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}\bar{T}}\big)$
subsections 3.2 and 3.3 $\ell^{2}\big(\mathcal{F}^{(1/2)}_{JJ}+\mathcal{F}^{(1/2)}_{\bar{J}\bar{J}}\big)$ $\ell^{4}\big(\mathcal{F}^{(1/2)}_{JJJJ}+\mathcal{F}^{(1/2)}_{\bar{J}\bar{J}\bar{J}\bar{J}}+\mathcal{F}^{(1/2)}_{TJJ}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{J}\bar{J}}$
$\mathcal{F}^{(1/2)}_{TJJ}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{J}\bar{J}}+\mathcal{F}^{(1/2)}_{TT}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}}+\mathcal{F}^{(1/2)}_{JJ}\mathcal{F}^{(1/2)}_{\bar{J}\bar{J}}\big)$
subsections 4.2 and 4.3 $\ell^{2}\mathcal{F}^{(1/2)}_{\varepsilon\varepsilon}$ $\ell^{4}\big(\mathcal{F}^{(1/2)}_{\varepsilon\varepsilon\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}_{T\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}_{\bar{T}\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}_{TT}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}}\big)$
Table 1: The nontrivial leading and subleading terms of fidelities in free massless boson and fermion theories. We need the universal contributions $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}}$ (2.30) with $\mathcal{X}=T,\bar{T},J,\bar{J},\varepsilon$, $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}\mathcal{X}}$ (2.36) with $\mathcal{X}=T,\bar{T}$, $\mathcal{F}^{(1/2)}_{T\mathcal{X}\mathcal{X}}$ (2.5) with $\mathcal{X}=J,\varepsilon$, and $\mathcal{F}^{(1/2)}_{\bar{T}\mathcal{X}\mathcal{X}}$ (2.5) with $\mathcal{X}=\bar{J},\varepsilon$, as well the specific contributions $\mathcal{F}^{(1/2)}_{JJJJ}$ (2.43), $\mathcal{F}^{(1/2)}_{\bar{J}\bar{J}\bar{J}\bar{J}}$ (2.44) and $\mathcal{F}^{(1/2)}_{\varepsilon\varepsilon\varepsilon\varepsilon}$ (2.48).
@@ 3.2 Case II
Using the contributions from current operators and stress tensor, $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}}$ (2.30) with $\mathcal{X}=J,\bar{J},T,\bar{T}$, $\mathcal{F}^{(1/2)}_{T\mathcal{X}\mathcal{X}}$ (2.5) with $\mathcal{X}=J$, $\mathcal{F}^{(1/2)}_{\bar{T}\mathcal{X}\mathcal{X}}$ (2.5) with $\mathcal{X}=\bar{J}$, $\mathcal{F}^{(1/2)}_{JJJJ}$ (2.43), and $\mathcal{F}^{(1/2)}_{\bar{J}\bar{J}\bar{J}\bar{J}}$ (2.44) , and organizing them in a factorized form
$F(\rho_{A},\sigma_{A})=\big[1+\ell^{2}\mathcal{F}^{(1/2)}_{JJ}+\ell^{4}\big(\mathcal{F}^{(1/2)}_{JJJJ}+\mathcal{F}^{(1/2)}_{TJJ}+\mathcal{F}^{(1/2)}_{TT}\big)+O(\ell^{6})\big]$
@@
we reproduce the short interval expansion of the fidelity between vertex operator states. The factorization reflects the holomorphic-antiholomorphic separation in the free boson theory.